• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh

    Log entry time 12:43:15 on December16,2007

    Entry number 221476

    keyword=Controlled access finished

    Right hrs NMR meter was replaced and that fixed the NMR problem. Thank you Ed and Jack fixing the problem. Asked for beam for normal production runs (target boiling test).